OpenStudy (anonymous):

An egg is dropped from the top of a fictitious tower, which stands 1,111 feet tall. determine when the egg will hit the ground and splatter.

OpenStudy (anonymous):

@Rachella

OpenStudy (anonymous):

ohh my god @perl is typing

OpenStudy (perl):

there is an equation from physics we can use y = 1/2 * g * t^2 1111 = 1/2 * 32 * t^2